Horizon Value of Free Cash Flows Current and projected free cash flows for Radell Global Operations are shown below. Actual 2016 2017 Projected 2018 2019 Free cash flow $606.82 $667.50 $707.55 $750.00 (millions of dollars) Growth is expected to be constant after 2018, and the weighted average cost of capital is 10.4%. What is the horizon (continuing) value at 2019 if growth from 2018 remains constant? Round your answer to the nearest dollar. Do not round intermediate calculations.?
Answer #1

Growth rate =(FCF in 2019 -FCF in 2018)/FCF 2018 =(750-707.55)/705.55 =6.0166%
Horizon value or Terminal Value =FCF in 2019*(1+growth)/(WACC-growth) =750*(1+6.0166%)/(10.4%-6.0166%)
=18139.4465 or 18139